# The Eisenbud–Goto conjecture for projectively normal varieties with mild singularities

### v1

@@ Abstract.
For a nondegenerate projective variety $X$, the Eisenbud–Goto conjecture asserts that $\operatorname{reg}X\leq\operatorname{deg}X-\operatorname{codim}X+1$. Despite the existence of counterexamples, identifying the classes of varieties for which the conjecture holds remains a major open problem. In this paper, we prove that the Eisenbud–Goto conjecture holds for $2$-very ample projectively normal varieties with mild singularities.
@@ 1. Introduction
In this paper, we prove that the Eisenbud–Goto conjecture holds for 2-very ample projectively normal varieties with mild singularities. Specifically, we prove the following.
@@
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with mild singularities in the following sense:
  1. (1)
    $X$ has factorial rational singularities; and
  2. (2)
    $\dim T_{p}X\leq\dim X+1$ for any $p\in X$.
Then
@@ 4. The birational double point formula
Here, $E$ is exceptional for $g$ means that the image of every irreducible component of $E$ has dimension lower than $E$. Intuitively, the birational double point formula decomposes the non-isomorphic locus of a certain birational morphism into two parts: the folding locus and the exceptional locus. These loci are denoted by $D$ and $E$ in the lemma respectively, and $D$ is called the double point divisor. The positivity of the double point divisor has been investigated in several works [3, 19, 35, 8].
@@ 5. The Eisenbud–Goto regularity bound
for the general point $y\in X$ or $\pi(x)$ is the vertex point of $X_{x_{1}}$. The latter case cannot occur since $\dim T_{\pi(x)}X_{x_{1}}\leq n+1$ as $\pi^{\prime}:X\setminus\{x_{1}\}\to X_{x_{1}}$ is isomorphic at $x$.
@@
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with mild singularities in the following sense:
  1. (1)
    $X$ has factorial rational singularities; and
  2. (2)
    $\dim T_{p}X\leq\dim X+1$ for any $p\in X$.
Then
@@
As before, denote $e:=\operatorname{codim}X$ and $d:=\deg X$. Since any locally factorial Cohen–Macaulay variety is Gorenstein (cf. [6, Corollary 3.3.19]), $X$ is Gorenstein. Hence by Theorem 5.6, $\mathcal{O}_{X}$ is (de)(d-e)-regular. From the long exact sequence

### v2

@@ Abstract.
For a nondegenerate projective variety $X$, the Eisenbud–Goto conjecture asserts that $\operatorname{reg}X\leq\operatorname{deg}X-\operatorname{codim}X+1$. Despite the existence of counterexamples, identifying the classes of varieties for which the conjecture holds remains a major open problem. In this paper, we prove that the Eisenbud–Goto conjecture holds for $2$-very ample projectively normal varieties with factorial, rational, hypersurface singularities and isolated Gorenstein singularities.
@@ 1. Introduction
In this paper, we prove that the Eisenbud–Goto conjecture holds for 2-very ample projectively normal varieties with mild singularities. Regarding mild singularities, we allow two types of singularities: one is the factorial, rational, hypersurface singularity and the other is the isolated Gorenstein singularity. Specifically, we prove the following.
@@
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with factorial, rational, hypersurface singularities and isolated Gorenstein singularities. Then
@@ 4. The birational double point formula
Here, $E$ is exceptional for $g$ means that the image of every irreducible component of $E$ has dimension lower than $E$. The divisor $D$ is called the double point divisor. The positivity of the double point divisor has been investigated in several works [3, 19, 35, 8].
@@ 5. The Eisenbud–Goto regularity bound
for the general point $y\in X$ or $\pi(x)$ is a vertex point of $X_{x_{1}}$. The latter case cannot occur since $\dim T_{\pi(x)}X_{x_{1}}\leq n+1$ as $\pi^{\prime}:X\setminus\{x_{1}\}\to X_{x_{1}}$ is isomorphic at $x$.
@@
Let $X\subseteq\mathbb{P}^{r}$ be a nondegenerate $2$-very ample projectively normal variety with factorial, rational, hypersurface singularities and isolated Gorenstein singularities. Then
@@
As before, denote $e:=\operatorname{codim}X$ and $d:=\deg X$. Since any factorial Cohen–Macaulay point is a Gorenstein point (cf. [6, Corollary 3.3.19]), $X$ is Gorenstein. Hence by Theorem 5.6, $\mathcal{O}_{X}$ is (de)(d-e)-regular. From the long exact sequence